Solve using the square root property
(x+8)^2=81

Answer by madhan_math(34)

Given (x+8)^2=81
(Using the square root property)
taking square root on both side, we get,
(x+8)=square root (81)
x+8 =9 (9*9=81)
Therefore, x=9-8
Hence x=1